# TraceLens: Question-Driven Debugging for Taint Flow Understanding

### v1

@@ Abstract
Inspired by the insight that sensemaking tool-generated results can be significantly improved by a QA inquiry process, we propose TraceLens, a first end-user question-answer style debugging interface for taint analysis. It enables a user to ask why, why-not, and what-if questions to investigate the existence of suspicious flows, the non-existence of expected flows, and the global impact of third-party library models. TraceLens performs speculative what-if analysis, to help a user in debugging how different connectivity assumptions affect overall results. A user study with 12 participants shows that participants using TraceLens achieved 21% higher accuracy on average, compared to CodeQL. They also reported a 45% reduction in mental demand (NASA-TLX) and rated higher confidence in identifying relevant flows using TraceLens. This shows TraceLens’s potential to significantly reduce sensemaking effort.
Figure 1: TraceLens supports interrogative debugging by enabling a user to ask why, why-not, and what-if questions about taint-analysis. A user can select one among templated queries (shown in Q) and contextualize their inquiry with respect to a specific source, sink, and 3rd party library model using a drop-down menu, shown in (D-SRC), (D-SNK), and (D-API). Once configured, a background question-and-answer analysis is conducted to help a user with their sensemaking process of taint analysis results. To aid in their sensemaking process about global connectivity, permissible and impermissible data flows, a user can see the result in a graph view (G) with color-code annotation.
I Introduction
@@
To prevent security vulnerabilities, developers use taint analysis. This technique tracks potentially dangerous data as it moves through a program. Data from untrusted sources, like user input, is marked as "tainted" The analysis then monitors how this tainted data spreads. If it reaches a critical point in the code, known as a "taint sink," a warning is generated. Tainted data can be made safe by "sanitizers," which are functions that clean or encode the data. For example, the java.net.URLEncoder.encode(String input, String encoding) function escapes potentially harmful characters, preventing security issues. Since analyzing the entire program is usually impossible, practical taint analysis requires configuring simplified models of external libraries [1]. These models abstract the behavior of third-party libraries, allowing an analysis to track taint flow without analyzing the libraries’ internals.
@@
Taint analysis tools often come with default configurations, including pre-built models of external libraries. However, these default models, often automatically generated, can be inaccurate. This is because they rely on assumptions, which may be incorrect, about how libraries handle data flow. Incorrect configurations can lead to significant problems. For example, if a model incorrectly allows tainted data to flow through a library function that should sanitize it, the analysis will produce unexpected taint flows (i.e, false positives) [2, 3]. This means it will report potential vulnerabilities that do not actually exist, creating extra, unexpected warnings for the user. Conversely, if a model incorrectly blocks the flow of tainted data when it should not, the analysis will produce missing flows (i.e., false negatives). This means it will miss real vulnerabilities, failing to report flows expected by the user.
@@
We introduce TraceLens, a novel end-user debugger that brings interactive, question-based sensemaking to taint analysis. Drawing inspiration from interrogative debugging principles, TraceLens enables developers to ask why, why not, and what if questions about their analysis configuration. TraceLens provides six customizable question templates, which are then concretized by a user. A user can make sense of currently detected taint flows and hypothetical flows through interactive QA. Unlike state-of-the-art tools like CodeQL that rely on list views, TraceLens visualizes these reported and hypothetical flows graphically, offering a more intuitive understanding of configuration changes. Below, we discuss two key features,Inquiry-based Sensemaking and Visualization of Global Impact in detail.
@@
In the Query Options pane of TraceLens (shown in Figure 1), users can select from pre-defined question templates. These templates allow users to investigate taint analysis results by asking specific questions about dataflows. To make these questions concrete, users specify the configuration they are interested in, including the source of the data, the destination (sink), and the models of third-party libraries involved.
@@
Visualization of Global Impact. Once a user initiates a query within TraceLens, specifying their target source, sink, and any relevant external API calls, TraceLens generates an interactive graph visualization, providing a holistic view of the taint flow. This graph, as seen in the G pane of Figure 1, visually maps the program’s taint flows as interconnected nodes, each representing a distinct stage in the taint flow. This visualization reveals the global impact of configuration choices through the network of connected nodes. To facilitate sensemaking, TraceLens employs a color-coded scheme: source nodes are highlighted in green, sink nodes in red, and external library nodes in orange. This color differentiation draws attention to the impact of configuration choices. Furthermore, templated queries are designed to illuminate how which third-party library’s model can impact multiple sinks, multiple sources, etc, making global impact salient through visualization.
User study. We conducted a within-subject study with a factorial crossover on 12 participants (graduate students and professional developers), who inspected taint-analysis warnings generated by CodeQL [4, 5]. Each participant answered eight questions per task (16 total) designed to reflect a realistic sensemaking process of tool-generated warnings. During these tasks, a user had to answer questions about the impact of third-party library models on taint analysis results, by identifying pass-through API calls, APIs that serve as sanitizers, etc. Participants increased the average question completion rate from 71% with the baseline CodeQL visualizer to 92% with TraceLens, especially on questions that require global reasoning of multiple taint flows. When using TraceLens, users were more accurate in identifying multiple sinks affected by the same third-party library model. For example, 50% of answers prepared with TraceLens are correct, whereas only 17% with the baseline.
We measured cognitive load via NASA-TLX questions. We found that TraceLens led to improvement on the participants self-reported mental demand, stress, and success rate. Considering the technology acceptance model [6], participants rated TraceLens higher on both Confidence (mean 4.3 vs. 2.1) and Ease of Use (4.3 vs. 1.9) than the baseline CodeQL. Qualitative feedback highlighted that color-coded graphs and dedicated template queries helped alleviate analysis paralysis [7]. Overall, these results underscore that TraceLens substantially enhances users’ ability to make sense of complex taint paths and reduce their mental workload through QA-based end-user debugging.
@@
  1. 1.
    Configuring taint analysis is tricky, often leading to unexpected results. To help users understand and fix these issues, TraceLens is the first interactive end-user debugger that enables the investigation of how different user configuration choices in modeling third-party libraries impact the taint analysis results.
  2. 2.
    TraceLens is equipped with template queries for ‘why’, ‘why not’, and ‘what if’ questions, which are automatically translated into logic queries; it enables users to concretize the template queries with concrete code names embedded in tool-generated warnings. TraceLens’s graphical interface eases the examination of the global impact of models, which are difficult to reason when viewing warnings in a list view, packaged with an existing taint analyzer.
  3. 3.
    We conduct a within-subject user study with a factorial crossover design (12 participants) to assess TraceLens’s effectiveness in sensemaking CodeQL-generated taint analysis warnings [4, 5]. TraceLens raises the average question completion rate from 71% to 92% and reduces mental demand (NASA-TLX) from 5.9 to 3.3, demonstrating clear improvement in the end-user debugging of taint analysis.
  4. 4.
@@
The remainder of this paper is organized as follows. Section II introduces a motivating example and the design goals of TraceLens. Section III presents TraceLens’s approach. Section IV provides the study design and results from our within-subject user study. Section V reports the results we obtained from our user study. Section VI discusses implications of our findings, and possible threats to validity. Section VII presents related work. Finally, we draw the conclusions of our work in Section VIII.
II Motivation
When using modern static taint analyzers such as CodeQL [5], developers struggle to identify why unexpected flows appear or why certain expected flows are missing. These tools rely on models of external libraries, which may not always match the developer’s expectations. This paper is concerned with the problem of end-user debugging of taint analysis—i.e., understanding the impact of user configuration choices. Two key end-user debugging challenges prevent developers to easily make sense of unexpected results.
@@ II Motivation
Prior debugging introduced WhyLine [8] to let users ask “why did” or “why didn’t” on a program trace, where the context of debugging is on runtime events (e.g., runtime errors). However, “why and why-not” questions remain largely unaddressed in the context of modern taint analysis tools. Developers want customizability over taint analysis [9], visual outputs of warnings and code [10], and the ability to inspect intermediate pass through nodes of the analysis [8]. Yet, typical dataflow or taint analysis interfaces do not systematically address higher-level inquiries such as “Why does data from source X reach sink Y?” or “Why is no warning raised for a known bug?
Table I: Template queries for why, why-not and what-if questions and corresponding English interpretation and logic queries.
@@
Table I shows six kinds of templated queries. In the following paragraphs, we detail the motivating scenario behind each query.
1. WhyFlow: Why is there a taint flow from a source X to a sink Y? Suppose that Alice sees an unexpected taint-analysis warning in CodeQL from a source user.getSSN() (untrusted data) to a sink log(SSN) (potential vulnerability) [9, 10, 8]. She suspects the culprit is an imprecise third-party library model; however, manually tracing a long chain of calls is overwhelming. The WhyFlow query highlights in one view the taint flow path from a source in green to a sink in red, displaying pass-through third-party API calls in orange, shown in Figure 2. Alice then notices an intermediate API call is encrypt(SSN). She then discovers that this API is a sanitizer, and thus this warning should have not reported and encrypt()’s model should be debugged.
Figure 2: WhyFlow: “Why is there a taint flow from a source to a sink?”
@@
Figure 3: WhyNotFlow: “Why is there no taint flow from a source to a sink?”
3. AffectedSinks: If we alter a third party library $Z$’s model, which sinks are affected? Alice wants to reason about the global impact of updating a third party library model [11]. Suppose that she sees an unexpected warning and considers marking a third-party library’s API as a sanitizer. However, she is concerned this update might suppress other warnings. Without TraceLens, she would need to sift through all reported warnings, manually trace each flow from the source, and check which sinks might become unreachable after her change—a time-consuming process. Figure 4 shows how AffectedSinks automates this what-if analysis, immediately revealing all red sinks that would be “killed.”.
Figure 4: AffectedSinks: “If we alter a third party library’s model, which sinks are affected?
@@
Suppose that Alice would like to know whether a known vulnerability reaching multiple sinks could be fixed at once [11, 12]. In a typical taint analyzer, it can be cumbersome to trace how a single piece of sensitive data such as an SSN propagates through multiple taint flows. Consequently, developers may not prefer to identify a common interception point instead of fixing one sink at a time. With DivergentSinks query, she can quickly locate a common point from the source that “splits” into multiple sinks. (see Figure 5).
@@
Figure 5: DivergentSinks: “Which third party library model could influence multiple taint flows from the same source?”
@@
To illustrate another scenario, imagine a single sensitive destination (sink) that receives data from multiple untrusted sources. Using the DivergentSources query, Alice can readily identify if these diverse sources converge towards the same vulnerable point. This allows her to pinpoint the specific model responsible for the convergence, i.e., the ‘culprit model’. Unlike traditional list-based warning views, TraceLens provides a visual representation of where these multiple taint paths intersect, which makes pinpointing the culprit model significantly easier.
@@
6. GlobalImpact: Which third party library’s model could have the largest global influence on multiple flows from a source to a sink? When multiple models could explain a false or missing warning, developers often prefer a conservative fix that impacts the fewest flows [13]. Without TraceLens, identifying the frequency of each API in all taint paths would require her to painstakingly review every single warning and count occurrences manually. GlobalImpact automatically computes how often each API appears across multiple paths, ranking them by frequency. In Figure 6, because extractParts(standardizedSSN) appears in more paths, it has a higher score, signaling a larger global impact.
@@
Figure 6: GlobalImpact: “Which third party library model could have the largest global influence on dataflows from a source to a sink?”
III Speculative Analysis for Inquiry-based Debugging
@@
Figure 7: A user selects a WhyNotFlow template query to investigate a suspected missing flow msg:HttpRequest to ErrorLogger.warn(). TraceLens concretizes a corresponding logic query. For instance, to explain a potential missing flow from msg: HttpRequest (node 1) to ErrorLogger.warn() (node 2), a plausible path via an intermediate node (node 3) is first identified by the rule whynot_node_pairs_answer(1,2,3). The query result is shown in the Graph View.
@@
TraceLens enables inquiry-based debugging for sensemaking to debug spurious flows or missing flows. As presented in Figure 7, we execute both the CodeQL taint query as well as a less restricted version of this query (Sections III-A), convert the taint analysis results into Soufflé facts and store them in a database (Section  III-B). Subsequently, TraceLens queries the database to provide answers to the questions posed using TraceLens, according to the user’s choice (Section  III-C). The answers are presented as an interactive graph visualization. In this section, we present the components of TraceLens, and how each of these components help the overall workflow. Before we elaborate on our approach, we state several assumptions in our design.
@@
Assumption 1. We presume that a user of TraceLens is familiar with the subject program that she is inspecting, meaning that they may recognize that some flows are spurious, unexpected flows and some expected flows are missing, serving a starting point for end-user debugging. This way, running an interrogative debugger can help narrow down the third-party library’s model after selecting a relevant template query for ‘why’ or ‘why-not’ questions.
@@
Assumption 4. The goal of TraceLens is not to frame ‘why’ and ‘why-not’ question asking as a classification problem of false positive or false negative. In fact, TraceLens never removes or adds any taint warnings reported by CodeQL. Rather, TraceLens helps users with end-user debugging—how their configuration choices such as source/sink/sanitizer definitions and third-party models may lead to spurious or missing paths. Given that the target users need to maintain and update third-party models, TraceLens aims to simplify this process.
III-A CodeQL
CodeQL is an open-source static analysis framework, developed and maintained by GitHub. Users can use a specialized Domain-Specific Language (DSL) to detect potential security vulnerabilities, code quality issues, or other custom patterns [14]. While CodeQL offers code scanning capabilities out of the box, its extendability via user-defined queries and models is particularly noteworthy.
A key factor in CodeQL’s precision and comprehensiveness lies in its third party library modeling. Each library API (including third-party dependencies) can be modeled to reflect how data flows through its methods. Recognizing that manual creation of such models is time-consuming and error-prone, the CodeQL team has explored automated and machine-learning-based techniques to infer library behaviors [9]. Nonetheless, incomplete or incorrect modeling remains a practical challenge, especially in large ecosystems where libraries evolve frequently without thorough documentation of their input-output contracts.

### v2

# WhyFlow: Interrogative Debugger for Sensemaking Taint Analysis

@@ Abstract.
Inspired by the insight that sensemaking tool-generated results can be significantly improved by a QA inquiry process, we propose WhyFlow, a first end-user question-answer style debugging interface for taint analysis. It enables a user to ask why, why-not, and what-if questions to investigate the existence of suspicious flows, the non-existence of expected flows, and the global impact of third-party library models. WhyFlow performs speculative what-if analysis, to help a user in debugging how different connectivity assumptions affect overall results. A user study with 12 participants shows that participants using WhyFlow achieved 21% higher accuracy on average, compared to CodeQL. They also reported a 45% reduction in mental demand (NASA-TLX) and rated higher confidence in identifying relevant flows using WhyFlow. This shows WhyFlow’s potential to significantly reduce sensemaking effort.
copyright: acmlicensedjournalyear: 2026conference: IEEE/ACM International Conference on Software Engineering; April 15–17, 2026; Rio de Janeiro, Brazil
1. Introduction
@@
To prevent security vulnerabilities, developers use taint analysis. This technique tracks potentially dangerous data as it moves through a program. Data from untrusted sources, like user input, is marked as "tainted" The analysis then monitors how this tainted data spreads. If it reaches a critical point in the code, known as a "taint sink," a warning is generated. Tainted data can be made safe by "sanitizers," which are functions that clean or encode the data. For example, the java.net.URLEncoder.encode(String input, String encoding) function escapes potentially harmful characters, preventing security issues. Since analyzing the entire program is usually impossible, practical taint analysis requires configuring simplified models of external libraries (Chibotaru2019, ). These models abstract the behavior of third-party libraries, allowing an analysis to track taint flow without analyzing the libraries’ internals.
@@
Taint analysis tools often come with default configurations, including pre-built models of external libraries. However, these default models, often automatically generated, can be inaccurate. This is because they rely on assumptions, which may be incorrect, about how libraries handle data flow. Incorrect configurations can lead to significant problems. For example, if a model incorrectly allows tainted data to flow through a library function that should sanitize it, the analysis will produce unexpected taint flows (i.e, false positives) (Livshits2009, ; Banerjee2023, ). This means it will report potential vulnerabilities that do not actually exist, creating extra, unexpected warnings for the user. Conversely, if a model incorrectly blocks the flow of tainted data when it should not, the analysis will produce missing flows (i.e., false negatives). This means it will miss real vulnerabilities, failing to report flows expected by the user.
@@
We introduce WhyFlow, a novel end-user debugger that brings interactive, question-based sensemaking to taint analysis. Drawing inspiration from interrogative debugging principles, WhyFlow enables developers to ask why, why not, and what if questions about their analysis configuration. WhyFlow provides six customizable question templates, which are then concretized by a user. A user can make sense of currently detected taint flows and hypothetical flows through interactive QA. Unlike state-of-the-art tools like CodeQL that rely on list views, WhyFlow visualizes these reported and hypothetical flows graphically, offering a more intuitive understanding of configuration changes. Below, we discuss two key features,Inquiry-based Sensemaking and Visualization of Global Impact in detail.
@@
In the Query Options pane of WhyFlow (shown in Figure 1), users can select from pre-defined question templates. These templates allow users to investigate taint analysis results by asking specific questions about dataflows. To make these questions concrete, users specify the configuration they are interested in, including the source of the data, the destination (sink), and the models of third-party libraries involved.
@@
Visualization of Global Impact. Once a user initiates a query within WhyFlow, specifying their target source, sink, and any relevant external API calls, WhyFlow generates an interactive graph visualization, providing a holistic view of the taint flow. This graph, as seen in the G pane of Figure 1, visually maps the program’s taint flows as interconnected nodes, each representing a distinct stage in the taint flow. This visualization reveals the global impact of configuration choices through the network of connected nodes. To facilitate sensemaking, WhyFlow employs a color-coded scheme: source nodes are highlighted in green, sink nodes in red, and external library nodes in orange. This color differentiation draws attention to the impact of configuration choices. Furthermore, templated queries are designed to illuminate how which third-party library’s model can impact multiple sinks, multiple sources, etc, making global impact salient through visualization.
User study. We conducted a within-subject study with a factorial crossover on 12 participants (graduate students and professional developers), who inspected taint-analysis warnings generated by CodeQL (avgustinov2016ql, ; szabo2023incrementalizing, ). Each participant answered eight questions per task (16 total) designed to reflect a realistic sensemaking process of tool-generated warnings. During these tasks, a user had to answer questions about the impact of third-party library models on taint analysis results, by identifying pass-through API calls, APIs that serve as sanitizers, etc. Participants increased the average question completion rate from 71% with the baseline CodeQL visualizer to 92% with WhyFlow, especially on questions that require global reasoning of multiple taint flows. When using WhyFlow, users were more accurate in identifying multiple sinks affected by the same third-party library model. For example, 50% of answers prepared with WhyFlow are correct, whereas only 17% with the baseline.
We measured cognitive load via NASA-TLX questions. We found that WhyFlow led to improvement on the participants self-reported mental demand, stress, and success rate. Considering the technology acceptance model (lee2003technology, ), participants rated WhyFlow higher on both Confidence (mean 4.3 vs. 2.1) and Ease of Use (4.3 vs. 1.9) than the baseline CodeQL. Qualitative feedback highlighted that color-coded graphs and dedicated template queries helped alleviate analysis paralysis (schwartz2015paradox, ). Overall, these results underscore that WhyFlow substantially enhances users’ ability to make sense of complex taint paths and reduce their mental workload through QA-based end-user debugging.
@@
  1. (1)
    Configuring taint analysis is tricky, often leading to unexpected results. To help users understand and fix these issues, WhyFlow is the first interactive end-user debugger that enables the investigation of how different user configuration choices in modeling third-party libraries impact the taint analysis results.
  2. (2)
    WhyFlow is equipped with template queries for ‘why’, ‘why not’, and ‘what if’ questions, which are automatically translated into logic queries; it enables users to concretize the template queries with concrete code names embedded in tool-generated warnings. WhyFlow’s graphical interface eases the examination of the global impact of models, which are difficult to reason when viewing warnings in a list view, packaged with an existing taint analyzer.
  3. (3)
    We conduct a within-subject user study with a factorial crossover design (12 participants) to assess WhyFlow’s effectiveness in sensemaking CodeQL-generated taint analysis warnings (avgustinov2016ql, ; szabo2023incrementalizing, ). WhyFlow raises the average question completion rate from 71% to 92% and reduces mental demand (NASA-TLX) from 5.9 to 3.3, demonstrating clear improvement in the end-user debugging of taint analysis.
  4. (4)
@@
The remainder of this paper is organized as follows. Section 2 introduces a motivating example and the design goals of WhyFlow. Section 3 presents WhyFlow’s approach. Section 4 provides the study design and results from our within-subject user study. Section 5 reports the results we obtained from our user study. Section 6 discusses implications of our findings, and possible threats to validity. Section 7 presents related work. Finally, we draw the conclusions of our work in Section 8.
2. Motivation
When using modern static taint analyzers such as CodeQL (szabo2023incrementalizing, ), developers struggle to identify why unexpected flows appear or why certain expected flows are missing. These tools rely on models of external libraries, which may not always match the developer’s expectations. This paper is concerned with the problem of end-user debugging of taint analysis—i.e., understanding the impact of user configuration choices. Two key end-user debugging challenges prevent developers to easily make sense of unexpected results.
@@ 2. Motivation
Prior debugging introduced WhyLine (ko2004designing, ) to let users ask “why did” or “why didn’t” on a program trace, where the context of debugging is on runtime events (e.g., runtime errors). However, “why and why-not” questions remain largely unaddressed in the context of modern taint analysis tools. Developers want customizability over taint analysis (nachtigall2022large, ), visual outputs of warnings and code (johnson2013don, ), and the ability to inspect intermediate pass through nodes of the analysis (ko2004designing, ). Yet, typical dataflow or taint analysis interfaces do not systematically address higher-level inquiries such as “Why does data from source X reach sink Y?” or “Why is no warning raised for a known bug?
Table 1. Template queries for why, why-not and what-if questions and corresponding English interpretation and logic queries.
@@
Table 1 shows six kinds of templated queries. In the following paragraphs, we detail the motivating scenario behind each query.
1. WhyFlow: Why is there a taint flow from a source X to a sink Y? Suppose that Alice sees an unexpected taint-analysis warning in CodeQL from a source user.getSSN() (untrusted data) to a sink log(SSN) (potential vulnerability) (nachtigall2022large, ; johnson2013don, ; ko2004designing, ). She suspects the culprit is an imprecise third-party library model; however, manually tracing a long chain of calls is overwhelming. The WhyFlow query highlights in one view the taint flow path from a source in green to a sink in red, displaying pass-through third-party API calls in orange, shown in Figure 2. Alice then notices an intermediate API call is encrypt(SSN). She then discovers that this API is a sanitizer, and thus this warning should have not reported and encrypt()’s model should be debugged.
Figure 2. WhyFlow: “Why is there a taint flow from a source to a sink?”
@@
Figure 3. WhyNotFlow: “Why is there no taint flow from a source to a sink?”
3. AffectedSinks: If we alter a third party library $Z$’s model, which sinks are affected? Alice wants to reason about the global impact of updating a third party library model (smith2015questions, ). Suppose that she sees an unexpected warning and considers marking a third-party library’s API as a sanitizer. However, she is concerned this update might suppress other warnings. Without WhyFlow, she would need to sift through all reported warnings, manually trace each flow from the source, and check which sinks might become unreachable after her change—a time-consuming process. Figure 4 shows how AffectedSinks automates this what-if analysis, immediately revealing all red sinks that would be “killed.”.
Figure 4. AffectedSinks: “If we alter a third party library’s model, which sinks are affected?
@@
Suppose that Alice would like to know whether a known vulnerability reaching multiple sinks could be fixed at once (smith2015questions, ; piskachev2019codebase, ). In a typical taint analyzer, it can be cumbersome to trace how a single piece of sensitive data such as an SSN propagates through multiple taint flows. Consequently, developers may not prefer to identify a common interception point instead of fixing one sink at a time. With DivergentSinks query, she can quickly locate a common point from the source that “splits” into multiple sinks. (see Figure 5).
@@
Figure 5. DivergentSinks: “Which third party library model could influence multiple taint flows from the same source?”
@@
To illustrate another scenario, imagine a single sensitive destination (sink) that receives data from multiple untrusted sources. Using the DivergentSources query, Alice can readily identify if these diverse sources converge towards the same vulnerable point. This allows her to pinpoint the specific model responsible for the convergence, i.e., the ‘culprit model’. Unlike traditional list-based warning views, WhyFlow provides a visual representation of where these multiple taint paths intersect, which makes pinpointing the culprit model significantly easier.
@@
6. GlobalImpact: Which third party library’s model could have the largest global influence on multiple flows from a source to a sink? When multiple models could explain a false or missing warning, developers often prefer a conservative fix that impacts the fewest flows (murphy2013design, ). Without WhyFlow, identifying the frequency of each API in all taint paths would require her to painstakingly review every single warning and count occurrences manually. GlobalImpact automatically computes how often each API appears across multiple paths, ranking them by frequency. In Figure 6, because extractParts(standardizedSSN) appears in more paths, it has a higher score, signaling a larger global impact.
@@
Figure 6. GlobalImpact: “Which third party library model could have the largest global influence on dataflows from a source to a sink?”
3. Speculative Analysis for Inquiry-based Debugging
@@
Figure 7. A user selects a WhyNotFlow template query to investigate a suspected missing flow msg:HttpRequest to ErrorLogger.warn(). WhyFlow concretizes a corresponding logic query. For instance, to explain a potential missing flow from msg: HttpRequest (node 1) to ErrorLogger.warn() (node 2), a plausible path via an intermediate node (node 3) is first identified by the rule whynot_node_pairs_answer(1,2,3). The query result is shown in the Graph View.
@@
WhyFlow enables inquiry-based debugging for sensemaking to debug spurious flows or missing flows. As presented in Figure 7, we execute both the CodeQL taint query as well as a less restricted version of this query (Sections 3.1), convert the taint analysis results into Soufflé facts and store them in a database (Section 3.2). Subsequently, WhyFlow queries the database to provide answers to the questions posed using WhyFlow, according to the user’s choice (Section 3.3). The answers are presented as an interactive graph visualization. In this section, we present the components of WhyFlow, and how each of these components help the overall workflow. Before we elaborate on our approach, we state several assumptions in our design.
@@
Assumption 1. We presume that a user of WhyFlow is familiar with the subject program that she is inspecting, meaning that they may recognize that some flows are spurious, unexpected flows and some expected flows are missing, serving a starting point for end-user debugging. This way, running an interrogative debugger can help narrow down the third-party library’s model after selecting a relevant template query for ‘why’ or ‘why-not’ questions.
@@
Assumption 4. The goal of WhyFlow is not to frame ‘why’ and ‘why-not’ question asking as a classification problem of false positive or false negative. In fact, WhyFlow never removes or adds any taint warnings reported by CodeQL. Rather, WhyFlow helps users with end-user debugging—how their configuration choices such as source/sink/sanitizer definitions and third-party models may lead to spurious or missing paths. Given that the target users need to maintain and update third-party models, WhyFlow aims to simplify this process.
3.1. CodeQL
CodeQL is an open-source static analysis framework, developed and maintained by GitHub. Users can use a specialized Domain-Specific Language (DSL) to detect potential security vulnerabilities, code quality issues, or other custom patterns (codeql_docs, ). While CodeQL offers code scanning capabilities out of the box, its extendability via user-defined queries and models is particularly noteworthy.
A key factor in CodeQL’s precision and comprehensiveness lies in its third party library modeling. Each library API (including third-party dependencies) can be modeled to reflect how data flows through its methods. Recognizing that manual creation of such models is time-consuming and error-prone, the CodeQL team has explored automated and machine-learning-based techniques to infer library behaviors (nachtigall2022large, ). Nonetheless, incomplete or incorrect modeling remains a practical challenge, especially in large ecosystems where libraries evolve frequently without thorough documentation of their input-output contracts.